OpenStudy (anonymous):

The price of a cup of coffee was $2.75 yesterday. Today, the price fell to $2.40, Find the percentage decrease. (( Round your answer to the nearest tenth of a percent.))

OpenStudy (anonymous):

divide the difference by original price .35/2.45= 1428571429 then times by 100 1428571429*100 = 14.258.......... rounded 14%

jhonyy9 (jhonyy9):

2,75 ----- 100% 2,40 ----- x% ------------------ x = 2,40*100 /2,75 = 87,27 100%-87,30% = 12,7 % rounded 13%